OpenStudy (freemap):

It takes a boat 4 h to travel 75 km downstream. Let x represent the boat’s speed in still water, and let y represent the water’s current in kilometers per hour. The water is moving at 3.5 kmph. What is the speed of the boat?

OpenStudy (freemap):

18.75 is my answer

OpenStudy (welshfella):

net speed will be x - y this equals distance / time = 75/4 so we have x - 3.5 = 75/4

OpenStudy (welshfella):

18.75 is correct

OpenStudy (freemap):

Thanks

OpenStudy (welshfella):

yw